# Do Not Worry if You Do Not Have Data: Building Pretrained Language Models Using Translationese

### v2

@@ Abstract
In this paper, we explore the utility of Translationese as synthetic data created using machine translation for pre-training language models (LMs). Pre-training requires vast amounts of monolingual data, which is mostly unavailable for languages other than English. Recently, there has been a growing interest in using synthetic data to address this data scarcity. We take the case of English and Indic languages and translate web-crawled monolingual documents (clean) into the target language. Then, we train language models containing 28M and 85M parameters on this translationese data (synthetic). We show that their performance on downstream natural language understanding and generative tasks is only 3.56% poorer on NLU tasks and 1.51% on NLG tasks than LMs pre-trained on clean data. Further, we propose the use of lightweight TinyLMs pre-trained on clean data to filter synthetic data efficiently which significantly improves the performance of our models. We also find that LMs trained on synthetic data strongly benefit from extended pretraining on a tiny fraction (10%) of clean data. We release the data we collected and created as a part of this work, IndicMonoDoc, the largest collection of monolingual document-level corpora, which we hope will help bridge the gap between English and non-English performance for large language models.
@@ 1 Introduction
Large language models (LLMs)Brown et al. (2020); Workshop et al. (2022); Almazrouei et al. (2023); Lin et al. (2022) have been able to perform very well on downstream tasks like MMLU Hendrycks et al. (2021), Big-Bench Srivastava et al. (2022), etc, and have even started to reach human potential in many of these tasks. But this performance has very largely been credited to their scale and the vast amount of data that they have been fed. Most of these language models (LMs) perform well in languages like English where abundant data is available Kudugunta et al. (2023), but a vast majority of languages don’t have comparable data as compared to English. As a consequence, many LLMs, both monolingual and multilingual, involving these languages still show poor performance for various downstream tasks. For example, the largest open source multilingual model BLOOM Workshop et al. (2022) covers 46 natural languages spanning 9 language families, but the top 5 languages comprise 74.14% of the data. Despite the benefits of multilingualism Dabre et al. (2020), this skew in data still means that the low-resource languages will not perform well.
@@
Figure 1: Comparison of NLU performance between language models trained on clean data and synthetic translationese data in English, Hindi, and Gujarati shows that filtering synthetic text and extending training on a clean subset (10%) can bring the performance of models trained on synthetic data closer to LMs trained only on clean data extracted from the web.
@@
Fortunately synthetic data is an option and previous works such as, but not limited to, back-translation Sennrich et al. (2016a), sequence distillation Kim and Rush (2016), also known as forward translation, etc. have shown that synthetic data obtained using machine translation (MT) can supplement resource scarcity and can significantly enhance model performance Popović et al. (2020); Gala et al. (2023). However, to the best of our knowledge, there has been no work on showing the effectiveness of synthetic data for training LMs. Furthermore, the quality of synthetic data is also important, which many works take for granted. While there are options such as round-trip-translation Moon et al. (2020) or referenceless neural quality estimation (QE) Rei et al. (2021), they either involve twice the compute or a reasonably large model not available for most languages, and this might not be optimal to determine the quality of synthetic documents efficiently. We thus consider TinyLMs Eldan and Li (2023) as an efficient alternative, which have been shown to model documents by their fluent paragraph generation capabilities.
@@
In this paper, we focus on English and Indic languages and present a comprehensive study of the utility of synthetic, translationese Gellerstam (1986), monolingual data obtained using machine translation (MT) for training LMs. We propose a simple framework that involves training small language models, TinyLMs, on original web-crawled data (clean) and then using them to filter synthetic data. We then compare LMs trained on clean and synthetic data followed by fine-tuning on downstream tasks, where we observe that, unfiltered synthetic data LMs perform slightly poor compared to LMs trained on clean data, but after filtering, it matches its performance. We also show that tuning these synthetic data LMs on small clean data leads to further improvements.
@@
Figure 2: Overview of our approach to pre-train language models using translationese data. We leverage rich monolingual corpora in the src language and scarce corpora in the tgt language. Our method involves employing a pre-trained machine translation model to translate src to tgt. We then filter the resulting text using a TinyLM trained solely on clean tgt monolingual data for perplexity. The filtered synthetic data can be used to further pretrain larger language models.
@@
b. Demonstrating the efficacy of language models trained on filtered synthetic data across a range of downstream tasks in natural language understanding and generation.
c. A new document-level monolingual corpora (IndicMonoDoc) consisting of 39.5B tokens worth of monolingual clean document-level data spanning 22 scheduled languages and English. This dataset, which is 2 times larger than IndicCorpV2 Doddapaneni et al. (2023), will be made public alongside our models and code [^1] https://github.com/meetdoshi90/TinyLM/ [^2] https://huggingface.co/datasets/cfilt/IITB-MonoDoc.
@@ 2 Related Work
This paper focuses on creating, filtering, and utilizing synthetic data for training TinyLMs.
Monolingual Data: Previous efforts of collecting monolingual corpora for Indic languages include but not limited to the EMILLE/CIIL corpus McEnery et al. (2000), HindMonoCorp Bojar et al. (2014), Leipzig corpus Goldhahn et al. (2012), IndicCorpv1 Kakwani et al. (2020) and more recently IndicCorpv2 Doddapaneni et al. (2023) which shows significant increase in size compared to previous Indic corpora. Although the size of IndicCorpv2 is large, it is a sentence-level corpora that can be used to train NLU models but language models with their increase in size can be trained on longer contexts that require long sequences of documents. In this work, we build upon existing monolingual corpora for Indic languages and also show the efficacy of using synthetic data.
Synthetic Data Generation and Quality Estimation: Synthetic data is valuable for NLP tasks like using back translations Sennrich et al. (2016a), Edunov et al. (2018) to enhance Machine translation Marie et al. (2020), Bogoychev and Sennrich (2019), Ni et al. (2022), or for tasks such as native language identification Goldin et al. (2018). However, there’s a limited exploration of using synthetic data for pretraining LMs due to issues like hallucination Maynez et al. (2020), and ungrounded non-factual text Thorne et al. (2018). Machine translation can help mitigate these issues, but translation errors may still occur, impacting text quality. Evaluating synthetic text quality with round-trip-translation (RTT) BLEU scores requires twice the compute and is prone to errors. Alternative evaluation methods like BARTScore Yuan et al. (2021), T5Score Qin et al. (2023), MQM & COMET Rei et al. (2020) require large-scale models or human annotations, limiting scalability. Approaches like KenLM Heafield (2011) have been used to filter monolingual corpora based on perplexity. We take inspiration from this and show that efficient TinyLM with just 28M parameters can be used to filter synthetic text.
TinyLMs: Language models as small as 10M parameters have been shown to produce fluent and consistent stories with almost perfect grammar Eldan and Li (2023) which means LMs even at a small scale have language understanding. Challenges like BabyLM Warstadt et al. (2023) focus on improving LMs with a fixed data budget which enables exhaustive study of LM development methodologies, which can then be applied to larger LMs. We aim to utilize lightweight TinyLMs for efficient filtering of synthetic documents.
@@ 3 Methodology
In this section, we describe our framework for leveraging synthetic data for LM training. This consists of monolingual data curation from the web (clean), training a TinyLM with it, translation of clean data, using the aforementioned TinyLM to filter synthetic data, and then using this filtered data for training a larger LM to be used for downstream tasks. Our framework is described in Figure 2.
3.1 Monolingual Data
@@
Web Crawled (Clean): Following Doddapaneni et al. (2023); Rae et al. (2022); Team et al. (2022), for all languages of interest, we a. obtain a list of URLs to be crawled via word-level n-grams passed to a search engine, b. after URL deduplication, we crawl all applicable webpages, c. automatically and manually Ortiz Suárez et al. (2019); Abadji et al. (2022) filter out unwanted text like HTML tags and emoticons, d. use language detection-based (LID) filtering using cld3333https://github.com/google/cld3 and IndicLID-FTN model Madhani et al. (2023a) to discard languages not of interest, e. perform document filtering to remove offensive text using toxic words list provided by Team et al. (2022), f. merge all the filtered corpus with Wikipedia, OSCAR Ortiz Suárez et al. (2019) and some dumps of mC4 Xue et al. (2021) and finally, g. perform deduplication at paragraph level using Murmurhash algorithm [^4] https://pypi.org/project/mmh3/ with a 128-bit unsigned hash for each monolingual split of the corpora.
Translationese (Synthetic): We utilize state-of-the-art MT models like IndicTrans2 Gala et al. (2023) to generate translationese data. We use beam search with a beam value of 5 to translate English tokens from the aforementioned crawled corpus to the languages of interest. Most MT models have a maximum token limit, and thus we split the documents using Moses Sentence Splitter [^5] https://pypi.org/project/mosestokenizer/ to perform translations into the target language at the sentence level and then merge again to form documents. Our experiments also focus on synthetic English data translated from Hindi and Gujarati.
@@ 3.2 Tiny Language Models (TinyLMs)
TinyLMs are simply tiny versions of language models inspired by Eldan and Li (2023). We follow the Transformer architecture Vaswani et al. (2017) used by Eldan and Li (2023) and train it using only clean monolingual documents. Instead of using learned positional encodings, we use RoPE embeddings Su et al. (2023) for better extrapolation to longer documents. We rely on Chinchilla scaling laws Hoffmann et al. (2022) and use compute optimal word tokens to train our models.
@@ 3.3 Synthetic Data Filtering
We first train TinyLMs on crawled data and then use them to compute perplexities on our synthetic documents ($W\in$ $w_{1},w_{2},\ldots,w_{N}$) using the equation:
@@
We also use our TinyLM to repair sentences in documents that exceed the maximum length of the MT models, but this happens in only 0.002% of such cases. We provide more details of our approach in Appendix C.
3.4 Final LM and Downstream Task Training
Using the filtered synthetic corpora, we train our final LMs which are comparatively larger, and fine-tune them for natural language understanding (NLU) tasks such as IndicGLUE Kakwani et al. (2020), GLUE Wang et al. (2018) and generation (NLG) benchmarks such as IndicNLG Kumar et al. (2022), summarization tasks Nallapati et al. (2016), Chen et al. (2021) and machine translation benchmarks Team et al. (2022), Gala et al. (2023).
4 IndicMonoDoc
Following the monolingual data curation strategy in Section 3, we crawl data for English and 22 Indic languages. As a result, we end up with IndicMonoDoc, with 27.5 billion tokens of Indic documents and 12 billion tokens of English documents for a total of 39.5 billion tokens of clean monolingual data. This is the largest ever for Indic languages, even surpassing Doddapaneni et al. (2023) by 2 times. We use IndicMonoDoc for all clean parts of our experiments. We report additional details of IndicMonoDoc in Appendix D.
4.1 Analysis of Crawled Corpora
Figure 3 gives an overview of the comparison of IndicMonoDoc which is a document-level corpus with that of IndicCorpV2 which is a sentence-level corpus. It is important to note that we paid special attention to the low-resource languages.
4.2 Analysis of Synthetic Data
We use IndicMonoDoc for the clean part of our experiments and translate parts of it for the synthetic experiments. Figure LABEL:fig:_variance_plot shows the perplexity mean and variance scores for TinyLM across token positions in the documents. This shows that on unseen documents, TinyLM shows higher variance on English documents generated by translating Gujarati documents from IndicMonoDoc as compared to English clean and English synthetic generated from Hindi. This also gives reason for the deterioration in results in Table 3 due to Gujarati documents. Figure 4 shows the distribution of lengths of filtered documents by TinyLMs showing that they do not add any bias for shorter documents during filtering. Although we do not experiment with all these languages, we believe that IndicMonoDoc will be an invaluable resource for Indic LMs.
@@ 5 Experiments
In this section, we describe the training procedure and datasets used for different models mentioned in Section 3. We pretrain and fine-tune all of the mentioned models from scratch in both mono and bilingual settings using a Causal Language Modeling (CLM) objective on NLG tasks and use a Linear classification head for all classification tasks. We specify the sample of the dataset used for pretraining and finetuning for each model and see the different effects of using synthetic corpora for pretraining.
@@ 5.1 Pretraining Data Settings
syn-XX_yy-unfiltered: Denotes synthetic monolingual documents in XX language generated by using yy as source during translation.
@@
BI-XX-YY Prefix: Denotes bilingual models trained using an equal mixture of monolingual corpora in XX and YY languages. We append an _syn prefix to either XX or YY if a synthetic version of that language is employed in training and a -parallel/nonparallel tag to denote whether a parallel version of XX and YY are used or not.
Note, for each split we only use the number of tokens that are required to reach the point of optimality Hoffmann et al. (2022) by the language model. We mention other details in Appendix B.
@@ 5.2 Implementation and Training
TinyLMs: We use Pytorch Lightning [^7] https://lightning.ai/docs/pytorch/stable/ for our implementations and train TinyLMs as described in Section 3.2. We use hidden sizes of 768 and have two variants, one with 4 layers (mini) and one with 12 layers (base; same as GPT2-base) with 28M and 85M non-embedding parameters respectively. The mini models are trained on clean data with sequence lengths of 4096888We keep long sequence lengths to be able to handle long documents for filtering. (mini-4k) for filtering synthetic documents as described in Section 3.3. On the other hand, for our main pre-training and downstream fine-tuning experiments, we train mini and base models with sequence lengths of 1024 (mini-1k and base-1k). Following Hoffmann et al. (2022) we use 2.4 billion word tokens per language to compute optimal training of base models. Since Gujarati has only 900M tokens in our dataset, whenever Gujarati is involved, we train only the mini-1k model. For models involving English and Hindi, we train both mini and base models. Additional details are in Appendix B.

### v3

# Pretraining Language Models Using Translationese

@@ Abstract
In this paper, we explore the utility of Translationese as synthetic data created using machine translation for pre-training language models (LMs) for low-resource languages (LRLs). Our simple methodology consists of translating large amounts of web-crawled monolingual documents (clean) into the LRLs, followed by filtering the translated documents using tiny LMs trained on small but clean LRL data. Taking the case of Indian languages, we pre-train LMs from scratch with 28M and 85M parameters, and then fine-tune them for 5 downstream natural language understanding (NLU) and 4 generative (NLG) tasks. We observe that pre-training on filtered synthetic data leads to relative performance drops of only 0.87% for NLU and 2.35% for NLG, compared to pre-training on clean data, and this gap further diminishes upon the inclusion of a small amount of clean data. We also study the impact of synthetic data filtering and the choice of source language for synthetic data generation. Furthermore, evaluating continually pre-trained larger models like Gemma-2B and Llama-3-8B in few-shot settings, we observe that using synthetic data is competitive with using clean data. Our findings suggest that synthetic data shows promise for bridging the pre-training gap between English and LRLs.
Pretraining Language Models Using Translationese
@@ 1 Introduction
Large language models (LLMs) Brown et al. (2020); Luccioni et al. (2023); Almazrouei et al. (2023); Lin et al. (2022) have been able to perform very well on downstream tasks like MMLU Hendrycks et al. (2021), Big-Bench Srivastava et al. (2022), etc, and have even started to reach human potential in many of these tasks. But this performance has very largely been credited to their scale and the vast amount of data that they have been fed. Most of these language models (LMs) perform well in languages like English where abundant data is available Kudugunta et al. (2023), but a vast majority of languages don’t have comparable data as compared to English. As a consequence, many LLMs, both monolingual and multilingual, involving these languages still show poor performance for various downstream tasks. For example, the largest open source multilingual model BLOOM Luccioni et al. (2023) covers 46 natural languages spanning 9 language families, but the top 5 languages comprise 74.14% of the data. Even for models like mT5 (Xue et al., 2021), the top 10 of 107 languages account for more than 75.48% of the training data. Despite the benefits of multilingualism Dabre et al. (2020), this data skew still means that low-resource languages will underperform.
@@
Figure 1: Comparing NLU performance in English, Hindi, and Gujarati shows that filtering synthetic data and adding 10% clean data improves models, approaching the performance of those trained only on clean web data.
@@
Fortunately synthetic data is an option and previous works such as, but not limited to, back-translation Sennrich et al. (2016a), sequence distillation Kim and Rush (2016), also known as forward translation, etc. have shown that synthetic data obtained using machine translation (MT) can supplement resource scarcity and can significantly enhance model performance Popović et al. (2020); Gala et al. (2023). However, to the best of our knowledge, there has been no work on showing the effectiveness of synthetic data for pre-training LMs. Furthermore, the quality of synthetic data is also important, which many works take for granted. While round-trip-translation Moon et al. (2020) or referenceless neural quality estimation (QE) Rei et al. (2021) are viable, they either involve twice the compute or a reasonably large model not available for most languages, and this might not be optimal to determine the quality of synthetic documents efficiently. We thus consider TinyLMs Eldan and Li (2023) as an efficient alternative, which have been shown to model documents by their fluent paragraph generation capabilities.
@@
In this paper, we focus on Indic languages such as Hindi, Gujarati, and Marathi, and present a comprehensive study of the utility of synthetic monolingual data, also called translationese Gellerstam (1986), obtained using machine translation (MT) for pre-training LMs. We propose a simple framework that involves training tiny language models, henceforth TinyLMs, on original web-crawled data (clean) and then using them to filter synthetic data. We then compare LMs of different scales pre-trained on clean and synthetic data followed by fine-tuning on natural language understanding (NLP) and generation (NLG) downstream tasks, where we observe that, while unfiltered synthetic data based LMs are inferior compared to LMs trained on clean data, filtering leads to performance comparable to the latter. We further show that tuning these synthetic data LMs on small clean data leads to further improvements. We also show that these trends hold when continually pre-training LLMs such as Gemma-2B and Llama-3-8B.
@@
Figure 2: Overview of our approach to pre-train language models using translationese data. We leverage rich monolingual corpora in the src language and scarce corpora in the tgt language. Our method involves employing a pre-trained machine translation model to translate src to tgt. We then filter, using perplexity, the resulting text using a TinyLM trained solely on clean tgt monolingual data. The filtered synthetic data can be used to further pretrain larger language models.
@@
b. Demonstrating the efficacy of language models (up to Llama-3-8B) trained on filtered synthetic data across a range of NLU and NLG tasks for low resource Indic languages.
c. A new document-level monolingual corpora (IndicMonoDoc) consisting of 39.5B tokens worth of monolingual clean document-level data spanning 22 scheduled languages and English [^1] Our code and datasets are released at https://github.com/meetdoshi90/TranslationesePretraining.
@@ 2 Related Work
This paper focuses on creating, filtering, and utilizing synthetic data to train TinyLMs.
Monolingual Data: Previous efforts to collect monolingual corpora for Indic languages include the EMILLE/CIIL corpus McEnery et al. (2000), HindMonoCorp Bojar et al. (2014), Leipzig corpus Goldhahn et al. (2012), IndicCorpv1 Kakwani et al. (2020a), and IndicCorpv2 Doddapaneni et al. (2023). While IndicCorpv2 is large, it is sentence-level and suitable for NLU models but not for longer contexts needed by language models. We extend these corpora and demonstrate the effectiveness of using synthetic data.
Synthetic Data Generation and Quality Estimation: Synthetic data aids NLP tasks like back translation for machine translation Sennrich et al. (2016a); Edunov et al. (2018); Marie et al. (2020); Bogoychev and Sennrich (2019); Ni et al. (2022) and native language identification Goldin et al. (2018). However, using synthetic data for pretraining LMs is less explored due to hallucination Maynez et al. (2020) and ungrounded text Thorne et al. (2018). Evaluation methods like RTT BLEU scores are computationally intensive, while others like BARTScore Yuan et al. (2021), T5Score Qin et al. (2023), MQM, and COMET Rei et al. (2020) require large-scale models or human annotations, limiting scalability. Approaches like KenLM Heafield (2011) have been used to filter monolingual corpora based on perplexity.
Transfer Learning and Cross-Lingual Fine-Tuning: Approaches like translate-train, as described by Conneau et al. (2018), involve fine-tuning a multilingual PLM using machine-translated training data and evaluating in the target language. Oh et al. (2022) combined translate-train and translate-test for improved cross-lingual fine-tuning. In contrast, our work focuses on pretraining language models and exploring how synthetic text impacts pretraining and various downstream NLU and NLG tasks.
TinyLMs: Small LMs, even with 10M parameters, produce fluent and consistent text Eldan and Li (2023). Challenges like BabyLM Warstadt et al. (2023) focus on improving LMs within a fixed data budget. We take motivation from this and leverage TinyLMs for efficient filtering of synthetic documents.
@@ 3 Methodology
In this section, we describe our framework for leveraging synthetic data for LM training. This process consists of collecting monolingual (clean) data from the web for low-resource languages, training TinyLMs with it, translating clean data from a high resource language such as English into low-resource languages, using the aforementioned TinyLMs to filter synthetic data, and then using this filtered data to train LMs for downstream tasks. Our framework is illustrated in Figure 2.
@@
3.1 Collecting Clean Monolingual Corpora
Following Doddapaneni et al. (2023); Rae et al. (2022); Team et al. (2022), for all languages of interest, we a. obtain a list of URLs to be crawled via word-level n-grams passed to a search engine, b. after URL deduplication, we crawl all applicable webpages, c. automatically and manually Ortiz Suárez et al. (2019); Abadji et al. (2022) filter out unwanted text like HTML tags and emoticons, d. use language detection-based (LID) filtering using cld3222https://github.com/google/cld3 and IndicLID-FTN model Madhani et al. (2023a) to discard languages not of interest, e. perform document filtering to drop documents containing offensive text using toxic words list provided by Team et al. (2022), f. merge all the filtered corpus with Wikipedia, OSCAR Ortiz Suárez et al. (2019) and some dumps of mC4 Xue et al. (2021) and finally, g. perform deduplication at paragraph level using Murmurhash algorithm [^3] https://pypi.org/project/mmh3/ with a 128-bit unsigned hash for each monolingual split of the corpora.
We crawl data for English, with Indic context, and 22 Indic languages. As a result, we end up with IndicMonoDoc, with 27.5 billion tokens worth of Indic language documents and 12 billion tokens of English documents for a total of 39.5 billion tokens of clean monolingual data. This is larger than the corpora released by Doddapaneni et al. (2023), surpassing it by almost 2 times. We use IndicMonoDoc for all experiments with clean data. Figure 3 gives an overview of the comparison of IndicMonoDoc. Note that, creation of IndicMonoDoc is important since IndicCorpV2 is a sentence-level corpus, and training LMs need a document-level corpus. It is important to note that we paid special attention to the low-resource languages. In this paper we only use data corresponding to English, Hindi, Gujarati and Marathi. We report additional details of IndicMonoDoc in Appendix E.
3.2 Generating Translationese (Synthetic)
We utilize state-of-the-art MT models like IndicTrans2 Gala et al. (2023) to generate translationese data. Using beam search with a beam value of 5, we translate English tokens from the clean corpus to the target languages. Due to token limits in MT models, we split documents using the Moses Sentence Splitter [^4] https://pypi.org/project/mosestokenizer/ for sentence-level translations, then merge them back into documents. We use the 1B En-Indic version [^5] https://huggingface.co/ai4bharat/indictrans2-en-indic-1B of IndicTrans2 to translate 5B English tokens worth of documents from IndicMonoDoc into translationese data for Hindi, Marathi and Gujarati.
@@ 3.3 Tiny Language Models (TinyLMs)
TinyLMs are small language models inspired by Eldan and Li (2023). We use the Transformer architecture Vaswani et al. (2017) and train them with clean monolingual documents. RoPE embeddings Su et al. (2023) are used instead of learned positional encodings for handling long documents. Following Chinchilla scaling laws Hoffmann et al. (2022), we use compute-optimal word tokens. Although it is plausible to train a TinyLM on unfiltered translationese data to filter itself, our preliminary experiments revealed that they favor poor-quality data and hence we avoid this route.
@@ 3.4 Synthetic Data Filtering
We use these TinyLMs to filter the generated translationese data. We do this by using perplexity as a measure of document quality score. For language models, perplexity quantifies how well a model predicts a sequence of tokens. A lower perplexity indicates a natural document. It is calculated by:
@@
where the negative log-likelihood measures the error of the model’s predictions. While calculating perplexity over a sequence of tokens, $W\in$ $w_{1},w_{2},\ldots,w_{N}$ we skip the first $s$ tokens where $s=10$, $e=1024$ and calculate loss until only the first $e$ tokens of the document. We find setting $e$ to larger values can lead to higher variance in the document scores due to the size of the TinyLM. Following initial analyses, we choose $s$ and $e$ such that we remove the high uncertainty of the language at the start of an unseen document and avoid penalizing longer documents due to the fragility of the extrapolation ability of TinyLM [^6] During experiments we saw that these TinyLMs can only go up to a certain context length before deteriorating in quality.. Note that it is important to choose $e$ such that the language model gives a uniform estimate of perplexity over an already seen sequence of tokens $\in$ $w_{s},w_{s+1},\ldots,w_{e}$. For our experiments, we use the TinyLMs to score all synthetically generated translationese data and calculate a document score using the above method. Following, Laurençon et al. (2022) we do subsampling by thresholding document perplexity scores except Laurençon et al. (2022) did it using Ken-LM Heafield (2011) and we do it using our TinyLM. We keep the threshold value such that we include enough documents to reach the computed optimal token count for pretraining experiments.
@@ 4 Experiments
This section outlines the training procedures and datasets for the models described in Section 3. We pre-train decoder only LMs and fine-tune all models from scratch in monolingual and bilingual settings using the causal language modeling (CLM) objective for NLG tasks and a linear classification head for classification tasks. We specify the dataset samples used for pretraining and fine-tuning, and analyze the effects of synthetic corpora on pretraining.
@@ 4.1 Pretraining Data Settings
syn-XX_yy-unfiltered: Denotes synthetic monolingual documents in XX language generated by using yy as a source during translation.
@@
BI-XX-YY Prefix: Denotes bilingual models trained using an equal mixture of monolingual corpora in XX and YY languages. We append an _syn prefix to either XX or YY if a synthetic version of that language is employed in training, and a -parallel/nonparallel tag to denote whether a parallel version of XX and YY are used or not.
Note, for each split we only use the number of tokens that are required to reach the point of optimality Hoffmann et al. (2022) by the language model.
@@ 4.2 Implementation and Training
TinyLMs: We use Pytorch Lightning [^8] https://lightning.ai/docs/pytorch/stable/ for our implementations and train TinyLMs as described in Section 3.3 for filtering. We use hidden sizes of 768 and have two variants, one with 4 layers (mini) and one with 12 layers (base; same as GPT2-base) with 28M and 85M non-embedding parameters respectively. The mini models are trained on clean data with sequence lengths of 4096999We keep long sequence lengths to be able to handle long documents for filtering. (mini-4k) for filtering synthetic documents as described in Section 3.4. On the other hand, for our main pre-training and downstream fine-tuning experiments, we train mini and base models with sequence lengths of 1024 (mini-1k and base-1k). Following Hoffmann et al. (2022) we use 2.4 billion word tokens per language to compute optimal training of base models. Since Gujarati has only 900M tokens in our dataset, whenever Gujarati is involved as the target, we train only the mini-1k model. For models involving English and Hindi, we train both mini and base models. Additional details are in Appendix B.